OpenStudy (anonymous):

steve ran a distance of 150 meters in 1 1/2 minutes. what is his speed?

OpenStudy (anonymous):

100 mph or 6000 mph?

OpenStudy (jdoe0001):

is it required in mph?

OpenStudy (jdoe0001):

\(\bf \cfrac{150mt}{1\frac{1}{2}min}\implies \cfrac{150mt}{\frac{3}{2}min}\implies \cfrac{150mt}{1}\cdot \cfrac{2}{3}min\)